Notes: Each packet begins with a 14 byte Ethernet header A missing byte “__” denotes an...

Notes:
Each packet begins with a 14 byte Ethernet header
A missing byte “__” denotes an unknown or corrupted byte
Embedded protocol numbers in decimal are: TCP = 6; UDP = 17; and ICMP = 1

====================================================================================

Question 1

01 00 5e 00 00 02 00 0f f8 a0 c0 80 08 00 45 00

00 30 00 00 00 00 01 11 c5 3d 0a 7b 0a 03 e0 00

00 02 07 c1 07 c1 00 1c 85 03 00 00 08 03 0a 64

0a 00 63 69 73 63 6f 00 00 00 0a 7b 0a 01

  1. What is the IP version and IP header length?   
  1. What is the embedded protocol?
  1. What is the embedded protocol payload size?

Question 2

78 2b cb 9d 39 f3 00 0f f8 a0 b1 c0 08 00 45 00

00 28 09 7b 40 00 f4 06 70 13 d0 55 28 50 0a 7b

0a 21 00 50 36 bd d0 83 65 86 aa 7b 52 fb 50 10

10 2c 27 d9 00 00

  1. What is the IP version and IP header length?
  1. What is the embedded protocol?   
  1. What is the embedded protocol payload size?

Question 3

78 2b cb 9d 39 f3 00 0f f8 a0 b1 c0 08 00 45 00

00 __ 09 7b 40 00 f4 06 70 13 d0 55 28 50 0a 7b

0a 21 00 50 36 bd d0 83 65 86 aa 7b 52 fb 50 10

10 2c 27 d9 00 00 00 00 0a 26 d6 1e 5f 3a 33 12

  1. What is the IP version and IP header length?   
  1. What is the embedded protocol?
  1. What is the embedded protocol payload size?   
Answer #1

Given, each packet starts with 14 byte ethernet header.

Q1.

(a) 1st 4 bits of the packet denote IP version. 15th byte is 0x45. Hence, first 4 bits i.e. 4 is the IP version. Next 4 bits denote header length in 32-bit words. i.e. 5 32-bit words are there in header length i.e. header is 20 bytes.

(b) 10th byte in the header gives embedded protocol. i.e. 0x11, which means UDP (decimal 17 equivalent to 0x11).

(c) From (a), header is 20 bytes, hence count all bytes and subtract 14 and 20, and that will be the embedded protocol payload size. Here, 28 bytes.

Q2. Steps are same as above.

(a)15th byte is 0x45. Hence, IP version is 4. Header length = 5 32-bit words = 20 bytes.

(b) 10th byte in header is 0x06. which means TCP.

(c) Embedded protocol payload size = 54-20-14 = 20 bytes.

Q3.

(a)15th byte is 0x45, i.e. IP version is 4. header length = 5 32-bit words = 20 bytes

(b) 10th byte in header is 0x06 which means TCP.

(c)Embedded protocol payload size = 64 - 20 - 14 = 30 bytes